OpenStudy (anonymous):

The inverse of f(x) = 2x - 10 is f raised to the negative one, of x equals one-half times x plus five

OpenStudy (anonymous):

true or false

OpenStudy (aum):

To find the inverse, do the following steps: 1. Replace f(x) with y 2. Replace x with y and y with x. 3. Solve for y 4. Replace y with \(f^{-1}(x)\)

OpenStudy (aum):

A second method you can use to verify an inverse is: If \(f(f^{-1}(x)) = x\) or \(f^{-1}(f(x)) = x\) then they are inverses of each other.

OpenStudy (anonymous):

im lost

OpenStudy (aum):

f(x) = 2x - 10 1. Replace f(x) with y: y = 2x - 10 2. Replace x with y and y with x: x = 2y - 10 3. Solve for y: x = 2y - 10 Add 10: x + 10 = 2y divide by 2: x/2 + 5 = y or y = 1/2*x + 5 4. Replace y with f−1(x): \(f^{-1}(x) = \frac 12 x + 5 \)

OpenStudy (anonymous):

oh ok